# Explicit quantum surrogates for quantum kernel models

### v1

@@ Abstract
Quantum machine learning (QML) leverages quantum states for data encoding, with key approaches being explicit models that use parameterized quantum circuits and implicit models that use quantum kernels. Implicit models often have lower training errors but face issues such as overfitting and high prediction costs, while explicit models can struggle with complex training and barren plateaus. We propose a quantum-classical hybrid algorithm to create an explicit quantum surrogate (EQS) for trained implicit models. This involves diagonalizing an observable from the implicit model and constructing a corresponding quantum circuit using an extended automatic quantum circuit encoding (AQCE) algorithm. The EQS framework reduces prediction costs, mitigates barren plateau issues, and combines the strengths of both QML approaches.
@@
Quantum machine learning has attracted significant interest in recent years as a potential application of quantum computers [1, 2]. Particularly interesting approaches to quantum machine learning are those based on the idea of a quantum feature map, which aims to gain performance improvement over classical techniques via the use of quantum states for storing features of given data. [3, 4, 5, 6, 7, 8, 9]. Models using quantum features can be broadly classified into two categories: the explicit quantum models based on parameterized quantum circuits [6, 7, 8] such as quantum circuit learning [9], and the implicit models based on quantum kernels [3, 4, 5]. The former makes predictions about an input from the output of a single quantum circuit, which first embeds the input into a quantum state and then applies a parameterized quantum circuit to select important features stored in the state. The model is therefore explicitly specified via the description of the circuit used for prediction. In contrast, the latter does so by calculating inner products of quantum features, that is, quantum kernel functions, among the data points and assessing them on a classical machine. It is therefore implicit in the sense that the quantum circuit itself does not describe the model. It has been demonstrated that both quantum models are capable of achieving good learning performance in benchmark tasks, as seen in numerical simulations [4, 10, 11, 12, 13, 14, 15] and experiments using real quantum hardware [3, 16, 17, 18, 14, 19, 20]. Furthermore, in artificially and carefully designed scenarios, these models have demonstrated rigorous advantages over traditional classical methodologies [21, 22, 23, 12, 24].
Both explicit and implicit models have their advantages and limitations. In the training phase, implicit models always achieve lower training errors than explicit models due to a fundamental result known as the representer theorem [25, 26]. However, implicit models can sometimes overfit on the training data, resulting in a decrease in generalization performance [27]. Explicit models can, in certain cases, avoid this overfitting problem [27]. However, explicit models not only have complex training landscapes but also involve additional difficulty in identifying a suitable structure of parameterized quantum circuits, i.e., ansatz. In addition, certain parameterized quantum circuits are known to encounter a trainability problem, i.e., the barren plateau (BP) issue [28, 29], which refers to exponentially vanishing gradients as the number of qubits used increases.
In the prediction phase, implicit models require the evaluation of quantum kernels $\mathcal{O}(M)$ times, where $M$ is the number of training data points. Assuming that $M$ is on the order of $10^{3}$-$10^{4}$, which is commonly applied to kernel methods, and given the current scarcity of access to quantum computing, these factors significantly reduce the applicability of implicit models. In contrast, explicit models only require the execution of a single quantum circuit for making a prediction. In short, there is a trade-off between implicit and explicit models, each with its pros and cons, and there is no framework that allows for the best of both worlds.
@@
In this work, we propose a quantum-classical hybrid algorithm to construct a quantum explicit model that acts as a surrogate for a trained quantum implicit model, which we call the explicit quantum surrogate (EQS), as shown in Fig. 1. Our algorithm is composed of two key steps. The first step is the scalable diagonalization of an observable determined by the trained implicit model. An important observation is that predictions made by implicit models can be regarded as the expectation values of observables given as a linear combination of quantum feature states over a training dataset of size $M$. It is therefore possible to diagonalize the observable with $\mathrm{poly}(M)$ resources, while the dimension of $n$-qubit quantum feature states is $2^{n}$. Furthermore, we find numerically that such observables are low-rank approximable and that considering $K\ll M$ eigenvectors is sufficient to achieve high prediction accuracy. The second step is the construction of a quantum circuit that approximately diagonalizes the observable. This allows us to evaluate the expectation value by measuring the resulting quantum states in the computational basis. For the construction of the circuit, we use an extended version of the automatic quantum circuit encoding (AQCE) algorithm, originally proposed by [30] for constructing a circuit that generates a target state without assuming any fixed ansatz structure. We extend AQCE to generate a circuit for an isometry to diagonalize a low-rank observable. This allows us to construct an explicit model that has a training loss almost equivalent to the trained implicit model. The generated EQS may provide us with insights about ansatz structures suitable for specific datasets. One can also expect to mitigate the BP phenomenon via EQS, as the EQS automatically finds the structure and does not depend on randomly initialized ansatzes, which induce BP.
@@
We evaluate the performance of EQS for classification tasks through numerical simulations. The prediction accuracy of EQS using the MNISQ dataset [31] (a 10-qubit quantum circuit dataset with $M=$10,000 data points) and the VQE-generated dataset [14] (a 12-qubit quantum circuit dataset with $M=$1,800 data points) decreases by less than 0.010 compared to trained implicit models. It is notable that these results are obtained by only considering $K\leq 10$ eigenvectors. Additionally, we find that, even when the isometry generated by AQCE has a fidelity of only 0.6, the decrease in prediction accuracy remains at the same level. This suggests that we do not need to optimize the circuit carefully, and the computational cost for optimizing the circuit construction step may be smaller than one might expect. Furthermore, we find that the gradients of the EQS obtained by our method are $10^{3}-10^{4}$ times larger than when initializing with random parameters, suggesting a significant relaxation of BP. From these observations, we believe that our proposal is a valuable tool not only for reducing the prediction cost of implicit models but also for potentially enhancing the trainability of explicit models.
@@
Preliminary.— Let us first define the notations and terms, which are mostly based on those used in Refs. [27, 5]. Let $\rho(\bm{x})=|\psi(\bm{x})\rangle\langle\psi(\bm{x})|=U(\bm{x})|0\rangle\langle 0|U^{\dagger}(\bm{x})$ be an $n$-qubit quantum feature state that encodes an input $\bm{x}$ generated via a feature-encoding quantum circuit $U(\bm{x})$. We define the explicit model as:
@@
Algorithm to construct explicit quantum surrogate.— Our algorithm consists of two main components. The first component is the eigenvalue decomposition of the observable $O_{\bm{\alpha},\mathcal{D}}$ of the trained implicit model. The second component is constructing a quantum circuit that produces these eigenvectors. Although there are many ways of achieving this, our choice is to employ the AQCE [30] extended for constructing isometry that allows us to generate circuit without assuming fixed ansatz. The overview of our proposed algorithm is shown in Fig. 1. We describe each component in the sequence below.
First, we perform training of the implicit model and diagonalize the observable $O_{\bm{\alpha},\mathcal{D}}$. As $\mathcal{S}=\operatorname{span}\{|\psi(\bm{x}_{m})\rangle\}$ is an invariant subspace of $O_{\bm{\alpha},\mathcal{D}}$, it is sufficient to diagonalize $O_{\bm{\alpha},\mathcal{D}}$ within this subspace. For example, we can employ the following process to achieve this. We first determine the set of orthogonal basis vectors $\{|e_{i}\rangle\}_{i=1}^{\operatorname{dim}(\mathcal{S})}$ of $\mathcal{S}$ using e.g. the Gram-Schmidt process. The inner products $\langle\psi(\bm{x}_{m})|\psi(\bm{x}_{m^{\prime}})\rangle$ required in the process can be obtained using the Hadamard test [32]. Next, we compute the matrix elements of $O_{\bm{\alpha},\mathcal{D}}$ with respect to the new basis $\{|e_{i}\rangle\}$. They can be calculated as
@@
We can then diagonalize it classically to obtain its eigenvalues $\{\lambda_{k}\}_{k=0}^{\operatorname{dim}(\mathcal{S})-1}$ and corresponding eigenvectors $\{|\lambda_{i}\rangle\}_{k=0}^{\operatorname{dim}(\mathcal{S})-1}$ expressed as linear combinations of |$\ket{\psi(\bm{x}_{m})}$. Using these, the implicit model can now be rewritten as $f_{\mathrm{implicit}}(\bm{x};\bm{\alpha})=\sum_{k=0}^{\operatorname{dim}(\mathcal{S})-1}\lambda_{k}\bra{\lambda_{k}}\rho(\bm{x})\ket{\lambda_{k}}.$ Note that it is often possible to truncate the sum at $K\ll M$ without significant performance decrease, as we will discuss later and show in the numerical experiments. Truncation retains important information in a model while reducing its complexity. It may offer advantages in mitigating overfitting, but we leave such an analysis as future work.
Next, we construct a quantum circuit $\mathcal{C}$ that satisfies $\mathcal{C}|k\rangle\approx|\lambda_{k}\rangle$ for $k=0,...,K-1$, where $|k\rangle$ denotes the computational basis states. For this purpose, we extend AQCE [30] to isometries. The original AQCE is an algorithm that generates a circuit that outputs a target state $|\Psi\rangle$ with the desired accuracy. It constructs quantum circuits by iteratively adding two-qubit unitary gates without assuming any fixed ansatz. The optimization is performed in a manner similar to those in tensor network methods [33]. A brief review of AQCE and the extensions done in this work are given in Appendix A and B, respectively. The implicit model can now be translated into an explicit model, that is, an EQS:
@@
where $O=\sum_{k=0}^{K-1}\lambda_{k}|k\rangle\langle k|$. It should be noted that the above process of generating EQS can be understood as quantum architecture search (QAS) for finding a well-performing circuit for explicit models [34]. We present an example of the quantum circuit structure found in our numerical simulations, described later in Appendix S4. The analysis of the quantum circuit structure found by EQS is an interesting direction to explore, but it is beyond the scope of this work and will be considered in future research.
The possibility of low-rank approximation significantly affects cost of the AQCE step. Even though it does not raise intrinsic exponential cost to the number of qubits within the iterations for circuit optimization, we would expect that the AQCE would become increasingly difficult when the number of vectors to be constructed, $K$, is large. The low-rankness can be guaranteed when $\operatorname{dim}(\mathcal{S})$ is small, which we argue here to be expected for quantum features |$\ket{\psi(\bm{x})}$ that are well-designed, in the sense that “similar” $\bm{x}$ are mapped to similar feature vectors |$\ket{\psi(\bm{x})}$. For supervised learning with input $\bm{x}$ and output $y$, we say data $\bm{x}$ and $\bm{x}^{\prime}$ are similar when corresponding $y$’s are equal or close. For such well-designed features, we can expect that a large portion of {|$\{\ket{\psi(\bm{x}_{m})}\}$ is linearly dependent, thus making $\operatorname{dim}(\mathcal{S})$ small. Such well-designedness also guarantees prediction performance. Ref. [22] (Eq. (8)) shows that, if we wish to predict $y$ in the form of $y=\mathrm{Tr}[A\rho(\bm{x})]$ for an unknown observable $A$ having a sufficiently small norm, $\mathrm{Tr}(A^{2})$, we can achieve a prediction error of $\mathcal{O}(\operatorname{dim}(\mathcal{S}))$. We therefore assume that for a well-designed feature that predicts well, the observable of the trained implicit model, $O_{\bm{\alpha},\mathcal{D}}$, admits a low-rank approximation.
@@
Numerical analysis of EQS performance.— To evaluate our algorithm, we derive EQS from the implicit model of the support vector machine (SVM) [35, 36] using quantum kernel [3, 4]. We then compare the classification accuracies using two datasets: the MNISQ dataset, [31], which contains 10,000 samples with 10 labels, and the VQE-generated dataset [14], which contains 1,800 samples with 6 labels. The MNISQ dataset is derived from the quantum encoding of the MNIST dataset [37] and consists of labeled quantum circuits with 10 qubits. The VQE-generated dataset, created using the VQE algorithm [38, 39], comprises labeled quantum circuits with 12 qubits and has the property that output states of circuits with different labels exhibit extremely low fidelity. They are provided in QASM format [40]. Circuits in these datasets are essentially labeled according to the similarity of the output states; those with the same label output similar quantum states. The QASM string is an input data $\bm{x}$ in this case, and we simply choose $U(\bm{x})$ as the circuit described by the string $\bm{x}$.
@@
To assess how low-rank approximations affect classification accuracy, we compare multiple EQSs, each with a different value of $K$. In addition, to evaluate how approximations arising within the use of our extended AQCE algorithm for generating isometries affect classification accuracy, we compare two different models. The first model is an EQS represented by Eq. (5) optimized by our algorithm until it satisfies $F^{(k)}:=\left|\left\langle k\left|\mathcal{C}^{\dagger}\right|\lambda_{k}\right\rangle\right|>0.6$ for all $k$. The second model, which we call the exact low-rank model, directly uses the results of low-rank approximations of $O_{\bm{\alpha},\mathcal{D}}$, thus avoiding the error induced by the AQCE step. We adopt the one-vs-rest strategy for multi-class classification; that is, we construct a multi-class classifier by combining multiple binary classifiers. Note that this requires us to construct $O_{\bm{\alpha},\mathcal{D}}$ and the corresponding circuit $\mathcal{C}$ defining EQS for each label $l$, which we denote by $O_{\bm{\alpha},\mathcal{D}}^{(l)}$ and $\mathcal{C}^{(l)}$. The accuracy shown in the results is defined as the number of correct classifications divided by the total number of test data points. The quantum circuit simulation is performed using Qulacs 0.5.6 [41] in a noiseless environment. We detail the experimental conditions in Appendix C and show the results in Fig. 2. The results for other datasets can be found in Appendix D, which are consistent with the results presented here.
@@
First, we discuss the behavior of the exact low-rank models. In Fig. 2 (a) and (b), we observe that their accuracy improves and approaches the accuracy of the original one as $K$ increases. For instance, with the MNISQ-MNIST dataset as shown in Fig. 2 (a), the exact low-rank model exhibits only a 0.010 decrease in accuracy compared to the implicit model at $K=10$. For the VQE-generated dataset in Fig. 2 (b), there is a mere 0.014 decrease in accuracy compared to the implicit model at $K=1$. This can be explained by the fact that |$\ket{\psi(\bm{x})}$ with different labels exhibit extremely low fidelity in this dataset. Therefore, looking at the fidelity between a state |$\ket{\psi(\bm{x})}$ for an unknown $\bm{x}$ and an average of |$\ket{\psi(\bm{x}_{m})}$ within the same label $l$ in a training set would be sufficient to classify $\bm{x}$. Indeed, the first eigenvector of $O_{\bm{\alpha},\mathcal{D}}^{(l)}$ has a high fidelity of over 0.7 with |$\ket{\psi(\bm{x})}$ belonging to the same label $l$ for most $l$’s, as shown in Fig. S3 (a) of Appendix E.
@@
Overall, results indicate that high accuracy can be achieved with $K\ll M$ and that low-rank approximation is effective, as $M$ is on the order of $10^{3}-10^{4}$ for each dataset. Low-rankness can also be seen from the mean value of the cumulative contribution ratio $\frac{\sum_{i=0}^{K-1}\lambda_{i}^{2}}{\sum_{i=0}^{M-1}\lambda_{i}^{2}}$, which is shown in Fig. S3 (b) of Appendix E. For the MNISQ-MNIST and VQE-generated datasets, the values are 0.798 at $K=10$ and 0.744 at $K=6$, respectively. We believe that such low-rank natures are reasonable because observables are constructed from linear combinations of the quantum features that encode training data, and the quantum features in this example are well-designed in a sense that data with the same label are mapped to similar quantum states.
@@
Next, we discuss the performance of the EQS models. Fig. 2 (a) and (b) show that the impact of approximations made in the AQCE step on accuracy is surprisingly small. The decrease in accuracy from the exact low-rank model is only $0.008$ for the MNIST dataset at $K=10$. No decrease is observed for the VQE-generated dataset at $K=6$. It should be emphasized that the condition imposed on fidelities is only $F^{(k)}>0.6$. We assume this is because the use of imperfect replication of the eigenvectors could change the decision boundaries established in the training phase but did not affect the prediction results due to the margin preserved by the SVM. Given that the currently available quantum computers are affected by a non-negligible amount of noise, the fact that low target fidelity results in minimal degradation of accuracy may offer an advantage when executing our method on actual devices.
Figure 3: The sum of squared gradients obtained by explicit models with EQS-initialized ansatz and randomly initialized ansatz. The vertical axis represents the sum of squared gradients at the first training step. The horizontal axis represents the target label classified by the explicit model.

### v2

@@ Abstract
Quantum machine learning (QML) leverages quantum states for data encoding, with key approaches being explicit models that use parameterized quantum circuits and implicit models that use quantum kernels. Implicit models often have lower training errors but face issues such as overfitting and high prediction costs, while explicit models can struggle with complex training and barren plateaus. We propose a quantum-classical hybrid algorithm to create an explicit quantum surrogate (EQS) for trained implicit models. This involves diagonalizing an observable from the implicit model and constructing a corresponding quantum circuit using an extended automatic quantum circuit encoding algorithm. The EQS framework reduces prediction costs, provides a powerful strategy to mitigate barren plateau issues, and combines the strengths of both QML approaches.
@@
Quantum machine learning (QML) is an ambitious field that seeks to enhance machine learning capabilities by leveraging the power of quantum computers [Biamonte2017-ht, Cerezo2022-xy, dunjko2018machine, schuld2021machine]. The ultimate goal is to demonstrate a “quantum advantage” by solving learning problems that are intractable for classical computers. A central and promising strategy toward this goal is the concept of a quantum feature map, which encodes data into high-dimensional quantum states [Havlicek2019-uo, Schuld2019-iz, Schuld2021-ma, Benedetti2019-ip, Cerezo2021-gk, Bharti2022-ie, Mitarai2018-rp]. By representing data in this quantum feature space, QML models aim to uncover complex patterns and correlations that are beyond the reach of classical techniques.
@@
Models using quantum features can be broadly classified into two categories: the explicit quantum models based on parameterized quantum circuits [Benedetti2019-ip, Cerezo2021-gk, Bharti2022-ie] such as quantum circuit learning [Mitarai2018-rp], and the implicit models based on quantum kernels [Havlicek2019-uo, Schuld2019-iz, Schuld2021-ma]. The former makes predictions about an input from the output of a single quantum circuit, which first embeds the input into a quantum state and then applies a parameterized quantum circuit to select important features stored in the state. The model is therefore explicitly specified via the description of the circuit used for prediction. In contrast, the latter does so by calculating inner products of quantum features, that is, quantum kernel functions, among the data points and assessing them on a classical machine. It is therefore implicit in the sense that the quantum circuit itself does not describe the model. It has been demonstrated that both quantum models are capable of achieving good learning performance in benchmark tasks, as seen in numerical simulations [Schuld2019-iz, Schuld2020-jb, Liu2018-wm, Jerbi2021-oo, Skolik2022-sj, Nakayama2023-zc, D4DD00098F] and experiments using real quantum hardware [Havlicek2019-uo, Zhu2019-ss, Peters2021-so, Haug2023-bb, Nakayama2023-zc, Kusumoto2021-kd, Bartkiewicz2020-hx]. Furthermore, in artificially and carefully designed scenarios, these models have demonstrated rigorous advantages over traditional classical methodologies [Liu2021-pe, Huang2021-to, Pirnay2022-si, Jerbi2021-oo, Dunjko2017-ra].
The choice between these two paradigms involves a nuanced trade-off. While implicit models can find an optimal solution to minimize empirical loss due to the Representer Theorem [Smola1998-zt, Scholkopf2002-mh], both approaches face fundamental challenges rooted in the curse of dimensionality [cerezo2023does]. For fidelity kernels, this can lead to poor generalization from exponential concentration [thanasilp2024exponential]. Conversely, explicit models face significant training challenges due to non-convex landscapes and the barren plateau (BP) phenomenon [McClean2018-sg, Cerezo2021-gg, Larocca_2025], as well as the difficulty of identifying a suitable ansatz. Although both concentration and BPs stem from the same root cause, their impact on trainability is distinct, as we detail in Appendix C. Indeed, a well-trained implicit model is not guaranteed to outperform an explicit one on all tasks [Jerbi2023-lq].
Acknowledging this complex landscape, our work addresses a specific and practical scenario. Since the “curse of dimensionality” renders any learning algorithm intractable on unstructured data, our work—like most successful machine learning—focuses on “learnable” problems where the data possesses sufficient structure for a global fidelity quantum kernel to be trained successfully. This success, however, reveals a critical bottleneck, i.e., the $O(M)$ prediction cost.
The central question our work addresses is therefore how we can systematically convert such a “successful-but-slow” implicit model into a “fast-and-practical” explicit model, preserving its high performance while achieving an $O(1)$ prediction cost.
@@
In this work, we propose a quantum-classical hybrid algorithm to construct a quantum explicit model that acts as a surrogate for a trained quantum implicit model, which we call the explicit quantum surrogate (EQS), as shown in Fig. 1. Our algorithm is composed of two key steps. The first step is the scalable diagonalization of an observable determined by the trained implicit model. An important observation is that this observable can be diagonalized with $\mathrm{poly}(M)$ resources, despite the $2^{n}$ dimension of the feature space, and we find numerically that it is often low-rank approximable. The second step is the construction of a quantum circuit that approximately diagonalizes the observable. For this purpose, we extend the automatic quantum circuit encoding (AQCE) algorithm [Shirakawa2021-uj]. Our extension enables the generation of a circuit $\mathcal{C}$ that creates an isometry for diagonalizing the low-rank observable, and importantly, does so without assuming any fixed ansatz structure. This constructed circuit allows us to evaluate the expectation value by simply measuring the resulting quantum states in the computational basis.
This two-step construction process, while powerful, constitutes the main source of the one-time resource cost for our framework. A key contribution of this work is therefore a comprehensive analysis of these required resources, including measurement shots, classical processing, and circuit depth scalability. As detailed in Appendix E, we show this cost is a justifiable investment for applications that require high prediction throughput, thanks to the $O(1)$ prediction efficiency of the resulting EQS.
This procedure allows us to construct an explicit model that has a training loss almost equivalent to the trained implicit model. Furthermore, the generated EQS provides insights about ansatz structures suitable for specific datasets and can be used as a high-quality starting point to mitigate the BP phenomenon, as it does not depend on a randomly initialized ansatz.
@@
We evaluate the performance of EQS for classification tasks through numerical simulations. The prediction accuracy of EQS using the MNISQ dataset [Placidi2023-kq] (a 10-qubit quantum circuit dataset with M=M=10,000 data points) and the VQE-generated dataset [Nakayama2023-zc] (a 12-qubit quantum circuit dataset with M=M=1,800 data points) decreases by less than 0.010 compared to trained implicit models. It is notable that these results are obtained by only considering $K\leq 10$ eigenvectors. Additionally, we find that, even when the isometry generated by AQCE has a fidelity of only $0.6$, the decrease in prediction accuracy remains at the same level. This suggests that we do not need to optimize the circuit carefully, and the computational cost for optimizing the circuit construction step may be smaller than one might expect. Furthermore, we demonstrate that our method effectively mitigates the BP phenomenon. A scaling analysis reveals that the EQS initialization suppresses the exponential decay of gradients observed in randomly initialized circuits, with the performance gap widening to more than 5 orders of magnitude at 16 qubits. From these observations, we believe that our proposal is a valuable tool not only for reducing the prediction cost of implicit models but also for potentially enhancing the trainability of explicit models. Our framework is positioned in contrast to alternative strategies, such as those that modify the kernel itself or create classical surrogates, as detailed in Appendix A.
@@
Preliminary.— Let us first define the notations and terms, which are mostly based on those used in Refs. [Jerbi2023-lq, Schuld2021-ma]. Let $\rho(\bm{x})=|\psi(\bm{x})\rangle\langle\psi(\bm{x})|=U(\bm{x})|0\rangle\langle 0|U^{\dagger}(\bm{x})$ be an $n$-qubit quantum feature state that encodes an input $\bm{x}$ generated via a feature-encoding quantum circuit $U(\bm{x})$. We define the explicit model as:
@@
Algorithm to construct explicit quantum surrogate.— Our algorithm consists of two main components. The first component is the eigenvalue decomposition of the observable $O_{\bm{\alpha},\mathcal{D}}$ of the trained implicit model. The second component is constructing a quantum circuit that produces these eigenvectors. Although there are many ways of achieving this, our choice is to employ the AQCE [Shirakawa2021-uj] extended for constructing an isometry that allows us to generate a circuit without assuming a fixed ansatz. The overview of our proposed algorithm is shown in Fig. 1. We describe each component in the sequence below.
First, we perform training of the implicit model and diagonalize the observable $O_{\bm{\alpha},\mathcal{D}}$. As $\mathcal{S}=\operatorname{span}\{|\psi(\bm{x}_{m})\rangle\}$ is an invariant subspace of $O_{\bm{\alpha},\mathcal{D}}$, it is sufficient to diagonalize $O_{\bm{\alpha},\mathcal{D}}$ within this subspace. For example, we can employ the following process to achieve this. We first determine the set of orthogonal basis vectors $\{|e_{i}\rangle\}_{i=1}^{\operatorname{dim}(\mathcal{S})}$ of $\mathcal{S}$ using e.g. the Gram-Schmidt process. The inner products $\langle\psi(\bm{x}_{m})|\psi(\bm{x}_{m^{\prime}})\rangle$ required in the process can be obtained using the Hadamard test [Schuld2018-eg]. Next, we compute the matrix elements of $O_{\bm{\alpha},\mathcal{D}}$ with respect to the new basis $\{|e_{i}\rangle\}$. They can be calculated as
@@
[…]
Next, we construct a quantum circuit $\mathcal{C}$ that satisfies $\mathcal{C}|k\rangle\approx|\lambda_{k}\rangle$ for k=0,,K1k=0,...,K-1, where $|k\rangle$ denotes the computational basis states. For this purpose, we extend AQCE to isometries. The original AQCE is an algorithm that generates a circuit that outputs a target state $|\Psi\rangle$ with the desired accuracy. It constructs quantum circuits by iteratively adding two-qubit unitary gates without assuming any fixed ansatz. The optimization is performed in a manner similar to that in tensor network methods [Evenbly2009-cq]. A brief review of AQCE and the extensions done in this work are given in Appendix F and G, respectively.
A key strength of our EQS framework is its robustness to the potential sub-optimality of the circuit construction step. While the AQCE algorithm is powerful, it must navigate a highly complex, non-convex optimization landscape, and thus provides no guarantee of finding a global optimum. Our numerical results demonstrate, however, that a perfect circuit is not required to build a high-performance EQS.
The implicit model can now be translated into an explicit model, that is, an EQS:
@@
where $O=\sum_{k=0}^{K-1}\lambda_{k}|k\rangle\langle k|$. This explicit formulation offers a significant advantage in prediction efficiency, not only in terms of circuit executions but also in statistical cost. While a prediction requires only a single type of circuit execution, a potential concern is the number of measurement shots needed for estimation. However, as we prove in Appendix E.2, the required sample complexity is determined by the properties of the observable $O$ and is independent of the system size $n$. This addresses a key challenge for the practical application of QML models and rigorously establishes the efficiency of the EQS prediction phase.
@@
It should be noted that the above process of generating EQS can be understood as quantum architecture search (QAS) for finding a well-performing circuit for explicit models [noauthor_undated-wm]. We present an example of the quantum circuit structure found in our numerical simulations, described later in Appendix S7. The analysis of the quantum circuit structure found by EQS is an interesting direction to explore, but it is beyond the scope of this work and will be considered in future research.
[…]
@@
Numerical analysis of EQS performance.— To evaluate our algorithm, we derive EQS from the implicit model of the support vector machine (SVM) [Pl1999, Wu2003-nv] using global fidelity quantum kernel [Havlicek2019-uo, Schuld2019-iz]. We then compare the classification accuracies using two datasets: the MNISQ dataset [Placidi2023-kq], which contains 10,000 samples with 10 labels, and the VQE-generated dataset [Nakayama2023-zc], which contains 1,800 samples with 6 labels. The MNISQ dataset is derived from the quantum encoding of the MNIST dataset [Lecun1998-dr] and consists of labeled quantum circuits with 10 qubits. The VQE-generated dataset, created using the VQE algorithm [Peruzzo2014-gz, Tilly2022-nh], comprises labeled quantum circuits with 12 qubits and has the property that output states of circuits with different labels exhibit extremely low fidelity. A detailed description of these datasets, including their origin and construction, is provided in Appendix H.1. They are provided in QASM format [Cross2022-pb]. Circuits in these datasets are essentially labeled according to the similarity of the output states; those with the same label output similar quantum states. The QASM string is an input data $\bm{x}$ in this case, and we simply choose $U(\bm{x})$ as the circuit described by the string $\bm{x}$.
@@
To assess how low-rank approximations affect classification accuracy, we compare multiple EQSs, each with a different value of $K$. In addition, to evaluate how approximations arising within the use of our extended AQCE algorithm for generating isometries affect classification accuracy, we compare two different models. The first model is an EQS represented by Eq. (5) optimized by our algorithm until it satisfies $F^{(k)}:=\left|\left\langle k\left|\mathcal{C}^{\dagger}\right|\lambda_{k}\right\rangle\right|>0.6$ for all $k$. The second model, which we call the exact low-rank model, directly uses the results of low-rank approximations of $O_{\bm{\alpha},\mathcal{D}}$, thus avoiding the error induced by the AQCE step. We adopt the one-vs-rest strategy for multi-class classification; that is, we construct a multi-class classifier by combining multiple binary classifiers. Note that this requires us to construct $O_{\bm{\alpha},\mathcal{D}}$ and the corresponding circuit $\mathcal{C}$ defining EQS for each label $l$, which we denote by $O_{\bm{\alpha},\mathcal{D}}^{(l)}$ and $\mathcal{C}^{(l)}$. The accuracy shown in the results is defined as the number of correct classifications divided by the total number of test data points. The quantum circuit simulation is performed using Qulacs 0.5.6 [Suzuki2021-nj] in a noiseless environment. While the main results are presented for a noiseless environment to clearly assess the core performance, a detailed analysis of the method’s robustness against realistic shot noise is provided in Appendix E. This analysis shows that the impact on both final accuracy and, crucially, on the construction cost is limited and manageable. We detail the experimental conditions in Appendix H.2 and show the results in Fig. 2. The results for other datasets can be found in Appendix I, which are consistent with the results presented here.
@@
[…]
@@
Overall, results indicate that high accuracy can be achieved with $K\ll M$ and that low-rank approximation is effective, as $M$ is on the order of $10^{3}-10^{4}$ for each dataset. The effectiveness of this approximation can also be seen from the mean value of the cumulative contribution ratio $\frac{\sum_{i=0}^{K-1}\lambda_{i}^{2}}{\sum_{i=0}^{M-1}\lambda_{i}^{2}}$, which is shown in Fig. S5 of Appendix J. For the MNISQ-MNIST and VQE-generated datasets, the values are 0.798 at $K=10$ and 0.744 at $K=6$, respectively. We believe that this high amenability to low-rank approximation is reasonable because observables are constructed from linear combinations of the quantum features that encode training data, and the quantum features in this example are well-designed in the sense that data with the same label are mapped to similar quantum states.
@@
Next, we discuss the performance of the EQS models. Fig. 2 (a) and (b) show that the impact of approximations made in the AQCE step on accuracy is surprisingly small. The decrease in accuracy from the exact low-rank model (which assumes a perfect isometry with fidelities of 1.0) is only $0.008$ for the MNIST dataset at $K=10$. No decrease is observed for the VQE-generated dataset at $K=6$. It should be emphasized that the condition imposed on fidelities is only $F^{(k)}>0.6$. This finding has significant practical implications for the resource cost of EQS construction. It reveals a favorable trade-off between circuit depth and prediction accuracy: by accepting a minor compromise in fidelity, which is sufficient for high accuracy, we can halt the resource-intensive AQCE algorithm much earlier. This directly translates to a significantly shallower circuit, mitigating one of the key construction costs detailed in Appendix E. We assume this is because the use of imperfect replication of the eigenvectors could change the decision boundaries established in the training phase, but did not affect the prediction results due to the margin preserved by the SVM. Given that the currently available quantum computers are affected by a non-negligible amount of noise, the fact that low target fidelity results in minimal degradation of accuracy may offer an advantage when executing our method on actual devices.
Figure 3: Median sum of squared gradients for explicit models with different initializations. For each target label, we compute the sum of squared gradients at the first training step. The horizontal axis indicates the number of qubits $n$. The vertical axis shows the median of these values across the target labels. For $n<16$, all six labels were used; for the $n=16$ point, a subset of four labels was used due to the high simulation cost.
To further validate the practical advantage of our approach, we also benchmarked EQS against simpler heuristic classifiers. These results, detailed in Appendix L, confirm that for complex tasks where the original kernel model excels, our EQS significantly outperforms these baselines, justifying the necessity of its more sophisticated construction.